OpenStudy (konradzuse):

integrate (sin(x))^3

OpenStudy (turingtest):

now you get to try to be a little clever

OpenStudy (turingtest):

why don't I let you try it first and see how you approach it?

OpenStudy (konradzuse):

\[\int\limits \sin^2(x) \sin(x)\] \[\frac{1}{2}\int\limits(1-\cos^2(x)) \sin(x)\]

OpenStudy (turingtest):

where did the 1/2 come from?

OpenStudy (turingtest):

I think you are mixing up trig identities

OpenStudy (konradzuse):

oh woops stupid lecture :P

OpenStudy (konradzuse):

1-cos(2x)/2?

OpenStudy (konradzuse):

it seems I could bring the 1/2 out in front?

OpenStudy (konradzuse):

half angle fomulas hmm and pytagorean formula... Seems you could use either...

OpenStudy (konradzuse):

I konfused myself :P

OpenStudy (turingtest):

but one choice makes for a two-step problem and the other choice makes for a 3-step problem

OpenStudy (konradzuse):

ok so lets make it easy....

OpenStudy (turingtest):

yes

OpenStudy (konradzuse):

\[\int\limits \sin(x)-\cos^2(x) \sin(x)dx :P\]

OpenStudy (konradzuse):

\[\int\limits \sin(x) dx- \int\limits \cos^2(x) \sin(x)dx\] u = cos(dx) du = -sin(x)dx \[\int\limits \sin(x) dx--\int\limits u^2du\]

OpenStudy (konradzuse):

\[-\cos(x) + \frac{u^3}{3} + c\]

OpenStudy (konradzuse):

\[-\cos(x) + \frac{\cos(x)^3}{3} + c\]

OpenStudy (konradzuse):

Ok it's right WOOT!

OpenStudy (anonymous):

Refer to the attachment.

OpenStudy (konradzuse):

hmm?

OpenStudy (konradzuse):

Hope you didn't do it all out all nice and put it into a tiff file... :P

OpenStudy (anonymous):

Part of a screen capture from WolframAlpha.